Motron has two bonds outstanding, Series E and Series F. Both bonds have face values of $10,000 and, because both bonds are backed by Motron, share a 5.00% YTM, The Series E is a zero coupon bond with a maturity in 5 years. The Series F, also maturing in 5 years, is a hybridized bond that pays no coupon for the first year; then pays S350 every six months for two years (four total payments); and finally makes four $850 payments in the last two years Series E Series F YTM Issuance Maturity Term Face Coupons/Yr Coupon 5.00% 2/1/2019 2/1/2024 YTM Issuance Maturity Term Face Coupons/Yr Coupon 1 Coupon 2 Coupon 3 5.00% 2/1/2019 2/1/2024 5 years 5 years $10,000 $10,000 Begins after year: 0 0% 0% annually 7% annually 17% annuali A. What is the current price of a Series E bond? (Hint: The convention for zero-coupon bonds is to calculate the return on a semi-annual basis.) O B о с. O D O E. O A. $6,717.35 6,851.82 7,127.83 7,811.98 7,835.26 B. What is the current price of a Series F bond? (Hint: Even if you utilize formulas to solve this, you should lay out and discount the stream of cash flows as a check on your work.) O B о с. O D O A.$10,098.06 11,034.54 11,822.58 12,163.81 12,497.18
Answer #1
value of zero coupon bond fave value/(1+r)^n   =10000/(1.025)^10 7811.98
Year cash flow present value of cash flow = cash flow/(1+r)^n r = 5/2 =2.5%
1 0 0
2 0 0
3 350 325.0097938
4 350 317.0827257
5 350 309.3490007
6 350 301.8039031
7 850 715.0754498
8 850 697.6345852
9 850 680.6191075
10 10850 8476.002659
value of bond sum of present value of cash flow 11822.58